The larger of two numbers is 8 more than the smaller number.
when the smaller number is multiplied by 2, the result is the same as when the larger number is increased by 4.

Answer by Theo(13342)

the larger of 2 numbers is 8 more than the smaller number.

let x = the larger number.
let y = the smaller number.

x = y + 8

when the smaller number is multiplied by 2, the result is the same as when the larger number is increased by 4.

2 * y = x + 4

you have 2 equations that need to be solved simultaneously.

they are:

x = y + 8
2 * y = x + 4

replace x in the second equation with its equivalent value of y + 8 from the first equation to get:

2 * y = (y + 8) + 4

simplify to get:

2 * y = y + 12

subtract y from both sides of the equation to get:

y = 12

since x = y + 8, then x = 20

2 * y = 2 * 12 = 24
x + 4 = 20 + 4 = 24

solution looks good.

x is equal to 20 and y is equal to 12.

the larger number is 8 more than the smaller number.

2 times the smaller number is equal to the larger number + 4.